STARK SPECTRUM SIMULATION OF X_{2}Y_{4} ASYMMETRIC MOLECULES: APPLICATION TO ETHY-LENE IN A MFI-TYPE HOST ZEOLITE

M. SANZHAROV, \underline{V. BOUDON}, M.LOËTE, N. ZVEREVA-LOËTE, A. BALLANDRAS, G. WE-BER, Laboratoire Interdisciplinaire Carnot de Bourgogne, UMR 5209 CNRS-Université de Bourgogne, 9 avenue Alain Savary, B.P. 47870, F-21078 Dijon Cedex, France; M. ROTGER, Groupe de Spectrométrie Moléculaire et Atmosphérique (GSMA), CNRS UMR 6089, Moulin de la Housse, BP 1039, Cases 16-17, F-51687 Reims Cedex 2, France.

We present the D_{2h}T DS − ST (D_{2h} − ST ARK Top Data System) program suite developed to simulate Stark spectra of any IR active rovibrational polyad of X_{2}Y_{4} (D_{2h}) asymmetric-top molecules. It is based on the D_{2h}T DS package, released for studying any rovibrational band or polyad in the abscence of an electric field ^{a}^{,}^{b}. The D_{2h}T DS − ST suite consists in a series of FORTRAN programs called by a script. For calculation of Stark spectra we obtained the expressions of the dipole moment and polarizability operators of X_{2}Y_{4} molecules, using a tensorial formalism ^{c} analogous to the one developed for tetrahedral and octahedral molecules ^{d}. The developed program suite was used to estimate an effective average field in a cavity of a MFI-type host zeolite by comparison, with vibrational absorption spectra of ethylene in MFI-zeolite that have been recorded.

aCh. Wenger, W. Raballand, M. Rotger and V. Boudon, J. Quant. Spectrosc. Radiat. Transfer 95 (2005) 521-538.

bCh. Wenger, V. Boudon, M. Rotger, J.-P. Champion and M. Sanzharov, J. Mol. Spectrosc. 251 (2005) 102-113.

cW. Raballand, M. Rotger, V. Boudon and M. Loëte, J. Mol. Spectrosc. 217 (2003) 239-248.

dV. Boudon, J.-P. Champion, T. Gabard, M. Loëte, F. Michelot, G. Pierre, M. Rotger, Ch. Wenger and M. Rey, J. Mol.

Spectrosc. 228 (2004) 620-634.